# Spatially–Structured Models of Viral Dynamics: A Scoping Review

### v1

@@ Abstract
There is growing recognition in both the experimental and modelling literature of the importance of spatial structure to the dynamics of viral infections in tissues. Aided by the evolution of computing power and motivated by recent biological insights, there has been an explosion of new, spatially–explicit models for within–host viral dynamics in recent years. This development has only been accelerated in the wake of the COVID–19 pandemic. Spatially-structured models offer improved biological realism and can account for dynamics which cannot be well-described by conventional, mean-field approaches. However, despite their growing popularity, spatially–structured models of viral dynamics are underused in biological applications. One major obstacle to the wider application of such models is the huge variety in approaches taken, with little consensus as to which features should be included and how they should be implemented for a given biological context. Previous reviews of the field have focused on specific modelling frameworks or on models for particular viral species. Here, we instead apply a scoping review approach to the literature of spatially–structured viral dynamics models as a whole to provide an exhaustive update of the state of the field. Our analysis is structured along two axes, methodology and viral species, in order to examine the breadth of techniques used and the requirements of different biological applications. We then discuss the contributions of mathematical and computational modelling to our understanding of key spatially-structured aspects of viral dynamics, and suggest key themes for future model development to improve robustness and biological utility.
@@ 1 Introduction
However, despite their advantages, spatial models have been underused in an application context. Spatially–structured models of viral dynamics are routinely cited far less frequently and are much less commonly used in biological applications compared to mean–field compartmental models. While this may be in part due to the simplicity of the latter approach (including the difficulty in calibrating spatial models to experimental data), this is also likely due to the variety of approaches taken to modelling spatial structure in viral infections, and a lack of consensus on what features should be included and how they should be implemented. Mean–field models are more or less united in the common language of ODEs (possibly with some slight modifications to account for, for example, age structure or stochasticity), yet spatially–structured models span a huge variety of methods: from PDEs to compartmental approaches, to agent–based approaches, or combinations thereof. To add to the complexity, the actual nature of the spatial structure of infections by different viruses in different host tissues is also widely varied, depending on whether the infection is in vivo or in vitro, whether the host cells are motile or packed in tissue, how permissive the tissue is to virion diffusion, and so on. There is therefore a need for a wide–reaching, intersectional analysis of the full breadth of spatially–structured viral dynamics models to make sense of the breadth of techniques used and the requirements of different applications.
@@
The remainder of this paper is structured as follows. In Section 2 we present the methodology used for our scoping review, including inclusion / exclusion criteria. We then present an overview of the key publications in Section 3, before moving on to our detailed review of the field. In Sections 4 and 5 we present the state of the field by methodology and virus species respectively, before discussing innately spatially dynamic models in Section 6. Finally, in Section 7, we present emerging trends in the future developments of spatial modeling of viral dynamics, before closing the paper in Section 8.
@@ 2.2 Inclusion and exclusion criteria
After merging results from both databases and screening for duplicates, we analysed the combined pool of studies and pruned for relevance. Our criterion for inclusion was studies that introduce or apply a spatially–explicit mathematical or computational model of within–host viral dynamics. We excluded articles that met one or more of the following criteria:
@@
We conducted three iterations of screening of the studies found by our search, applying the inclusion criterion to first the titles, then the abstracts, then the full text of the remaining studies, discarding those which clearly did not meet the inclusion criterion. We only applied the exclusion criteria at the full text level. Of the 3492 unique studies returned by our search, we initially found 92 that were relevant to this review. The largest subset of publications which were removed on the basis of the exclusion criteria above were articles which were deemed to have an analytical focus (we found 111).
After having applied our selection criteria, we then screened the references of each of the accepted articles to find any articles missed by our database search. This process found a further 84 relevant studies for a total of 176. We outline the general workflow of the literature search process in Figure 1. We conducted our search in English and only considered articles written in English. We included pre–prints in our search.
Figure 1: Workflow for the scoping review. Duplicates automatically detected in JabRef. Duplicates manually detected.
@@ 2.3 Categorisation
In order to analyse high–level trends in the literature, we collected for each accepted study the viral species modelled (where one was specified) and the model framework used. We considered continuum models, patch models, cellular automata, and multicellular models in this review, which we define as follows. Models in which the spatial domain and all model components are continuous in space (that is, they are represented by densities, rather than discrete agents) are classified as continuum models. These are overwhelmingly PDE models, especially reaction–diffusion systems. Of the models which include spatially–discrete regions, we define patch models as those which represent cells as a density. Such models typically comprise a lattice of connected spatial regions, each of which contains a (possibly stochastically–simulated) ODE system. Models which included only a few spatial regions which were hard–coded to represent specific organs or tissue regions were not included in our search. The remaining models to be classified comprise spatially–discrete regions and a discrete number of cells and can collectively be considered agent–based models. We subdivide these models into cellular automata and multicellular models based on whether the model considers a single spatial scale or multiple, respectively. For instance, a model in which the cells are taken to be discrete spatial regions and where the extracellular viral load is tracked as the quantity of viral density at each cell would be considered a cellular automaton, since the cell–scale is the only spatial scale considered by the model. By contrast, if extracellular viral load were represented as an overlying field governed by a PDE, for example, the model would be considered multicellular. Note that we consider models with a global virus reservoir which has equal access to all spatial regions as having a secondary spatial scale and therefore multicellular. We summarise this classification in Figure 2. We assign to each study a single modelling framework: for those which make use of multiple models of different frameworks, we categorise the article by the model with the highest level of spatial resolution used (according to our classification, which corresponds to the furthest to the right on the tree in Figure 2). For each study, we also recorded its number of citations according to Google Scholar as of September 2nd, 2024.
Figure 2: Classification for modelling frameworks discussed in this review. Note, the boundaries of the methodology classifications used here are maintained through out the reminder of this paper.
@@ 3 An overview of the use of spatially–structured models in viral dynamics
Our search procedure returned 176 publications, spanning 60 years of research across a diverse group of virus species and a wide array of methodologies. The final annotated list of publications retained for analysis can be found on FigShare [^1] https://doi.org/10.26188/27085984.v1. In Figure 3 we present a broad overview of the distribution of the studies found by our search over time, and by virus modelled and method used. Figures 3(a)–(b) shows the averaged number of publications per year over time, computed by taking the average number of publications per year across the five–year interval centred on a given year. For 2022 and 2023 we compute the average over a shorter interval (from two years prior, up to 2023). We do not show data from 2024. In Figure 3(a) we stratify these studies by virus species modelled and in Figure 3(b) we stratify by methodology. In Figure 3(c), we show the overall split of methodologies used and for which virus species.
@@
To our knowledge, the earliest spatial description of viral spread in tissue with mathematical models was by Koch in 1964, who used the diffusion (heat) equation to analyse the growth of viral plaques in vitro [115]. However, the first “model” in a more recognisable sense, and in some ways the birth of the field of spatially–structured modelling for viral dynamics, was proposed some decades later by Pandey in 1989 [157]. Pandey’s model used a cellular automaton approach for the dynamics of HIV, and most of the spatial viral dynamics models in the early 1990s followed in this vein, as can be seen in Figures 3(a)–(b) [157]. Figures 3(a)–(b) also show significant growth and diversification of the literature on spatial spread of viral infections in the first half of the 2000s, with the first spatially–structured models specifically for influenza, hepatitis B and C (HBV and HCV), oncolytic viruses and others, along with the first patch and multicellular models all published around a similar time. Since 2020 and the outbreak of the COVID–19 pandemic, there has been a spike in the development of spatial models for viral infections, with the average yearly number of publications roughly doubling from less than six in the mid–’10s to approximately 13 in 2023. Curiously, this cannot be entirely attributed to models for SARS–CoV–2, with models for the virus accounting for less than a third of recent publications (19 of 62 publications since 2020).
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
@@
Figure 3: Spatially–structured viral dynamics models over time (a) by virus species and (b) by methodology. (c) Overall split of models by methodology.
@@
As shown in Figure 3(c), continuum models formed the largest proportion of the studies found by our search with 56, although between cellular automaton (54) and multicellular (52) approaches, we observed that agent–based models in general were a more popular approach (106 publications combined). Early agent–based models were almost exclusively cellular automata, although this approach fell out of favour in the 2010s, where multicellular approaches rose to prominence, as can be seen in Figure 3(b). Patch models were the least popular approach with only 14 publications since 1985. These were mostly models for HIV and herpes simplex virus 2 (HSV–2) with six publications each, including the entire spatial modelling literature specifically focused on HSV–2. The other modelling approaches were also highly specific to particular viral species with the largest proportion of continuum models focused on oncolytic viruses and the largest share of cellular automaton models developed for HIV. Multicellular models, by contrast, were most commonly proposed for general viral infections and not a particular viral species.
@@
We sought to identify which of the publications found had been the most impactful. To account for the time since publication, we quantified impact by considering both a study’s overall number of citations, and the average number of citations it had received per year since publication, and imposed thresholds on both quantities. These constraints ensure that a publication has both reached a specified degree of influence overall, but also that it has been influential relative to the number of years since its publication. In Figure 4, we plot each publication found in our search against its overall number of publications and its average number of publications per year. Publications are colour–coded by virus and markers indicate methodology used. We opted for thresholds on the overall and per–year citation counts of 50 and 6, respectively, and found 25 studies that passed both thresholds. We show a timeline of these seminal publications in Figure 5. In the proceeding analysis, we make particular mention of these high–impact publications. In this review, articles that were included in our search are referenced alongside the authors’ names, and high–impact publications are indicated in bold.
@@
Figure 4: Selection of seminal papers. We show each accepted study for our scoping review by total number of citations and average citations per year since publication. Cutoffs for inclusion in our analysis are shown.
@@
Figure 5: Seminal spatially–structured viral dynamics models over time. Criteria for inclusion are more than 50 total citations and more than 6 citations per year on average since publication. Pre–print. Seminal publications are as follows: Bernaschi and Castiglione [32], Zorzenon dos Santos and Coutinho [224], Wu et al. [212], Wein et al. [205], Funk et al. [75], Beauchemin et al. [27], Friedman et al. [73], Xiao et al. [213], Mok et al. [146], Heilmann et al. (2010) [93], Mitchell et al. [145], Wodarz et al. [210], Yakimovich et al. [215], Heilmann et al. (2012) [94], Schiffer et al. [177], Bocharov et al. [38], Malinzi et al. [132], Friedman and Lai [72], Eriksen et al. [66], Imle et al. [99], Alzahrani et al. [9], Getz et al. [78], Quirouette et al. [166], Simmons et al. [190], and Sego et al. [183].
@@ 4.1 Continuum models
Figure 6: Continuum models summary. Seminal publications are as follows: Wu et al. [212], Wein et al. [205], Friedman et al. [73], Mok et al. [146], Bocharov et al. [38], Malinzi et al. [132], Friedman and Lai [72], Alzahrani et al. [9], and Quirouette et al. [166].
@@ 4.2 Patch models
Figure 7: Patch models summary. Seminal publications are as follows: Funk et al. [75], and Schiffer et al. [177].
@@
The patch approach can be considered an extension of compartmental ODE models. For example, some authors have considered distinct ODE systems for the upper and lower regions of the respiratory tract, where the two compartments are coupled by the flow of virions [112, 81]; others considered more complex networks of interconnected organ compartments [204] (note that we do not consider such models explicitly spatial and thus did not include these in our analysis). Patch models extend this approach to an arbitrary network or lattice of compartments. Among the patch models found by our search, a popular approach was to employ a regular structure to the patches, with either rectangular (Funk et al. [75], Byrne et al. [43], Zhang et al. [223]) or hexagonal packing of the patches (Schiffer et al. [179, 180, 177, 178, 176, 61]). The patch approach has the property that its spatial structure is only defined implicitly via the strength and network of the connectivity between the patches. As such, depending on the structure of the patch connectivity, such models can reflect complex biological topologies which are not easily described by continuum models, for example. Nakaoka et al. proposed a patch model for the spread of HIV infection between lymph nodes, explicitly representing the complex interconnectivity of the lymphoid tissue network, while treating the dynamics within each lymph node as well–mixed [151]. Cardozo et al. [48, 47] and Jagarapu et al. [105], by contrast, applied a patch model at the within–lymph node scale. Their model considered distinct lymphoid follicles, interconnected via a free blood and lymph compartment, where the spherical lymphoid follicles were in turn subdivided into radial shell patches [105, 48, 47].

### v2

@@ Abstract
There is growing recognition in both the experimental and modelling literature of the importance of spatial structure to the dynamics of viral infections within the host. Aided by the evolution of computing power and motivated by recent biological insights, there has been an explosion of new, spatially–explicit models for within–host viral dynamics in recent years. This development has only been accelerated in the wake of the COVID–19 pandemic. Spatially-structured models offer improved biological realism and can account for dynamics which cannot be well-described by conventional, mean-field approaches. However, despite their growing popularity, spatially–structured models of viral dynamics are underused in biological applications. One major obstacle to the wider application of such models is the huge variety in approaches taken, with little consensus as to which features should be included and how they should be implemented for a given biological context. Previous reviews of the field have focused on specific modelling frameworks or on models for particular viral species. Here, we instead apply a scoping review approach to the literature of spatially–structured viral dynamics models as a whole to provide an exhaustive update of the state of the field. Our analysis is structured along two axes, methodology and viral species, in order to examine the breadth of techniques used and the requirements of different biological applications. We then discuss the contributions of mathematical and computational modelling to our understanding of key spatially-structured aspects of viral dynamics, and suggest key themes for future model development to improve robustness and biological utility.
@@ 1 Introduction
However, despite their advantages, spatial models have been underused in an application context. Spatially–structured models of viral dynamics are routinely cited far less frequently and are much less commonly used in biological applications compared to mean–field compartmental models. While this may be in part due to the simplicity of the latter approach (including the difficulty in calibrating spatial models to experimental data), this is also likely due to the variety of approaches taken to modelling spatial structure in viral infections, and a lack of consensus on what features should be included and how they should be implemented. Mean–field models are more or less united in the common language of ODEs (possibly with some slight modifications to account for, for example, age structure or stochasticity), yet spatially–structured models span a huge variety of methods: from PDEs to compartmental approaches, to agent–based approaches, or combinations thereof. To add to the complexity, the actual nature of the spatial structure of infections by different viruses in different host tissues is also widely varied, depending on whether the infection is in vivo or in vitro, whether the host cells are motile or packed in tissue, how permissive the extracellular environment is to virion diffusion, and so on. There is therefore a need for a wide–reaching, intersectional analysis of the full breadth of spatially–structured viral dynamics models to make sense of the breadth of techniques used and the requirements of different applications.
@@
1.1 How to read this paper
The remainder of this paper is structured as follows. In Section 2 we present the methodology used for our scoping review, including the inclusion and exclusion criteria. We then present an overview of the key publications in Section 3, before moving on to our detailed review of the field. We do so by first examining the literature from the perspective of the methods and techniques used (in Section 4), and then from the perspective of the biological insights these models have generated (in Section 5). We discuss distinct topics in these two sections, and in instances where a publication is cited more than once, it is to discuss different aspects of that publication: for example, its mathematical implementation in one section, and its contribution to biological knowledge in another. These sections are structured by model type and viral species modelled, respectively, for easy reference. Next, in Section 6, we identify some of the main aspects of viral infection dynamics which have been identified in the literature as being inherently spatially structured, and discuss progress made in the modelling literature in describing and understanding these processes. Finally, in Section 7, we present emerging trends for future development of spatially–structured models of viral dynamics, before closing the paper in Section 8.
@@ 2.2 Inclusion and exclusion criteria
After merging results from both databases and screening for duplicates, we analysed the combined pool of studies and pruned for relevance. Our criterion for inclusion was studies that introduce or apply a spatially–explicit mathematical or computational model of within–host viral dynamics. While many models we found considered the viral infection of tissues within a single host, we also found a number of studies of interest which considered only in vitro systems, or bacterial hosts. We deemed these studies relevant and useful to our discussion, and therefore adopted a broad interpretation of “within–host viral dynamics” as the dynamics of viral infection at a between–cell scale. Given that neither “within–host” nor “between–cell” featured in our search string, this interpretation will not affect the studies included in the initial search.
We excluded articles that met one or more of the following criteria:
@@
We conducted three iterations of screening of the studies found by our search, applying the inclusion criterion to first the titles, then the abstracts, then the full text of the remaining studies, discarding those which clearly did not meet the inclusion criterion. We only applied the exclusion criteria at the full text level. Of the 3939 unique studies returned by our search, we initially found 93 that were relevant to this review. The largest subset of publications which were removed on the basis of the exclusion criteria above were articles which were deemed to have an analytical focus (we found 111).
After having applied our selection criteria, we then screened the references of each of the accepted articles to find any articles missed by our database search. This process found a further 84 relevant studies for a total of 177. We outline the general workflow of the literature search process in Figure 1. We conducted our search in English and only considered articles written in English. We included pre–prints in our search.
Figure 1: Workflow for the scoping review. We queried Scopus and Web of Science using our search string to find an initial list of potentially relevant studies, then sequentially refined the list of studies based on our inclusion and exclusion criteria. We then repeated this approach on the reference lists of the included studies to end up with our final list of accepted publications. Duplicates automatically detected in JabRef. Duplicates manually detected.
@@ 2.3 Categorisation
In order to analyse high–level trends in the literature, we collected for each accepted study the viral species modelled (where one was specified) and the model framework used. We considered continuum models, patch models, cellular automata, and multicellular models in this review, which we define as follows. Models in which the spatial domain and all model components are continuous in space (that is, they are represented by densities, rather than discrete agents) are classified as continuum models. These are overwhelmingly PDE models, especially reaction–diffusion systems. Of the models which include spatially–discrete regions, we define patch models as those which represent cells as a density. Such models typically comprise a lattice of connected spatial regions, each of which contains a (possibly stochastically–simulated) ODE system. Models which included only a few spatial regions which were hard–coded to represent specific organs or tissue regions were not included in our search. The remaining models to be classified comprise spatially–discrete regions and a discrete number of cells and can collectively be considered agent–based models. We subdivide these models into cellular automata and multicellular models based on whether the model considers a single spatial scale or multiple, respectively. For instance, a model in which the cells are taken to be discrete spatial regions and where the extracellular viral load is tracked as the quantity of viral density at each cell would be considered a cellular automaton, since the cell–scale is the only spatial scale considered by the model. By contrast, if extracellular viral load were represented as an overlying field governed by a PDE, for example, the model would be considered multicellular. Note that we consider models with a global virus reservoir which has equal access to all spatial regions as having a secondary spatial scale and therefore multicellular. We summarise this classification in Figure 2, and provide illustrations of how spatial structure is typically — but by no means exclusively — represented within each model framework. We assign to each study a single modelling framework: for those which make use of multiple models of different frameworks, we categorise the article by the model with the highest level of spatial resolution used (according to our classification, which corresponds to the bottom of the tree in Figure 2). For each study, we also recorded its number of citations according to Google Scholar as of September 2nd, 2024.
Figure 2: Classification for modelling frameworks discussed in this review. For each of the model frameworks discussed in this review (continuum models, patch models, cellular automata and multicellular models) we show an indicative diagram of how space is represented in the system. In continuum models, all quantities are modelled as densities in continuous space; in patch models, space is discretised into regions within which cells and other system quantities are represented as densities; in cellular automata, space is discretised into (typically smaller) regions, within which cells are represented by discrete counts. Commonly, as we depict here, discrete regions correspond to the cells themselves. Multicellular models, like cellular automata, include discretisation of space and track discrete counts of cells (again commonly discretising space into the regions occupied by the cells themselves) but also describe some other spatial scale, such as a continuous virus density.
@@ 3 An overview of the use of spatially–structured models in viral dynamics
Our search procedure returned 177 publications, spanning 60 years of research across a diverse group of virus species and a wide array of methodologies. The final annotated list of publications retained for analysis can be found on FigShare [^1] https://doi.org/10.26188/27085984.v2. In Figure 3 we present a broad overview of the distribution of the studies found by our search over time, and by virus modelled and method used. Figures 3(a)–(b) shows the averaged number of publications per year over time, computed by taking the average number of publications per year across the five–year interval centred on a given year. For 2022 and 2023 we compute the average over a shorter interval (from two years prior, up to 2023). We do not show data from 2024. In Figure 3(a) we stratify these studies by virus species modelled and in Figure 3(b) we stratify by methodology. In Figure 3(c), we show the overall split of methodologies used and for which virus species.
@@
To our knowledge, the earliest spatial description of viral spread in tissue with mathematical models was by Koch in 1964, who used the diffusion (heat) equation to analyse the growth of viral plaques in vitro [22]. However, the first “model” in a more recognisable sense, and in some ways the birth of the field of spatially–structured modelling for viral dynamics, was proposed some decades later by Pandey in 1989 [23]. Pandey’s model used a cellular automaton approach for the dynamics of HIV, and most of the spatial viral dynamics models in the early 1990s followed in this vein, as can be seen in Figures 3(a)–(b) [23]. Figures 3(a)–(b) also show significant growth and diversification of the literature on spatial spread of viral infections in the first half of the 2000s, with the first spatially–structured models specifically for influenza, hepatitis B and C (HBV and HCV), oncolytic viruses and others, along with the first patch and multicellular models all published around a similar time. Since 2020 and the outbreak of the COVID–19 pandemic, there has been a spike in the development of spatial models for viral infections, with the average yearly number of publications roughly doubling from less than six in the mid–’10s to approximately 13 in 2023. Curiously, this cannot be entirely attributed to models for SARS–CoV–2, with models for the virus accounting for less than a third of recent publications (19 of 63 publications since 2020).
@@
Figure 3: Spatially–structured viral dynamics models over time (a) by virus species and (b) by methodology. Here we plot the five–year moving average number of publications per year (for 2022 and 2023, we show the average number of publications from two years prior, up to 2023; we do not show data for 2024). (c) Overall split of models by methodology. The number of publications for each method is given in brackets.
@@
As shown in Figure 3(c), continuum models formed the largest proportion of the studies found by our search with 56, although between cellular automaton (54) and multicellular (53) approaches, we observed that agent–based models in general were a more popular approach (107 publications combined). Early agent–based models were almost exclusively cellular automata, although this approach fell out of favour in the 2010s, where multicellular approaches rose to prominence, as can be seen in Figure 3(b). Patch models were the least popular approach with only 14 publications since 1985. These were mostly models for HIV and herpes simplex virus 2 (HSV–2) with six publications each, including the entire spatial modelling literature specifically focused on HSV–2. The other modelling approaches were also highly specific to particular viral species with the largest proportion of continuum models focused on oncolytic viruses and the largest share of cellular automaton models developed for HIV. Multicellular models, by contrast, were most commonly proposed for general viral infections and not a particular viral species.
@@
We sought to identify which of the publications found had been the most impactful. Though clearly not a complete picture of a study’s impact within the field, we chose to quantify this using citation data in order to limit editorial bias in the selection. To account for the time since publication, we considered both a study’s overall number of citations, and the average number of citations it had received per year since publication, and imposed thresholds on both quantities. These constraints were intended to identify publications that have reached a specified degree of influence overall, but have also been influential relative to the number of years since publication. In Figure 4, we plot each publication found in our search against its overall number of publications and its average number of publications per year. Publications are colour–coded by virus and markers indicate methodology used. We opted for thresholds on the overall and per–year citation counts of 50 and 6, respectively, and found 25 studies that passed both thresholds. We show a timeline of these highly–cited publications in Figure 5. In the proceeding analysis, we make particular mention of these highlighted publications. In this review, articles that were included in our search are referenced alongside the authors’ names, and highly–cited publications are indicated in bold.
@@
Figure 4: Identification of papers as highly–cited. We show each accepted study for our scoping review by total number of citations and average citations per year since publication. Cutoffs for definition as highly–cited in our analysis are shown. Citation data was collected from Google Scholar on September 2nd, 2024.
@@
Figure 5: Highly–cited spatially–structured viral dynamics models over time. Criteria for inclusion are more than 50 total citations and more than 6 citations per year on average since publication. Pre–print. Highly–cited publications are as follows: Bernaschi and Castiglione [24], Zorzenon dos Santos and Coutinho [25], Wu et al. [26], Wein et al. [27], Funk et al. [28], Beauchemin et al. [29], Friedman et al. [30], Xiao et al. [31], Mok et al. [32], Heilmann et al. (2010) [33], Mitchell et al. [34], Wodarz et al. [35], Yakimovich et al. [5], Heilmann et al. (2012) [36], Schiffer et al. [37], Bocharov et al. [38], Malinzi et al. [39], Friedman and Lai [40], Eriksen et al. [41], Imle et al. [2], Alzahrani et al. [42], Getz et al. [9], Quirouette et al. [43], Simmons et al. [44], and Sego et al. [8].
@@ 4.1 Continuum models
Figure 6: Continuum models summary. Left: Average number of publications using a continuum model per year. Centre: Continuum models by viral species studied. Right: Highly–cited continuum model publications over time. Highly–cited publications are as follows: Wu et al. [26], Wein et al. [27], Friedman et al. [30], Mok et al. [32], Bocharov et al. [38], Malinzi et al. [39], Friedman and Lai [40], Alzahrani et al. [42], and Quirouette et al. [43].
@@ 4.2 Patch models
Figure 7: Patch models summary. Left: Average number of publications using a patch model per year. Centre: Patch models by viral species studied. Right: Highly–cited patch model publications over time. Highly–cited publications are as follows: Funk et al. [28], and Schiffer et al. [37].
@@
The patch approach can be considered an extension of compartmental ODE models. For example, some authors have considered distinct ODE systems for the upper and lower regions of the respiratory tract, where the two compartments are coupled by the flow of virions [82, 83]; others considered more complex networks of interconnected organ compartments [84] (note that we do not consider such models explicitly spatial and thus did not include these in our analysis). Patch models extend this approach to an arbitrary network or lattice of compartments. Among the patch models found by our search, a popular approach was to employ a regular structure to the patches, with either rectangular (Funk et al. [28], Byrne et al. [85], Zhang et al. [86]) or hexagonal packing of the patches (Schiffer et al. [87, 88, 37, 89, 90, 91]). The patch approach has the property that its spatial structure is only defined implicitly via the strength and network of the connectivity between the patches. As such, depending on the structure of the patch connectivity, such models can reflect complex biological topologies which are not easily described by continuum models, for example. Nakaoka et al. proposed a patch model for the spread of HIV infection between lymph nodes, explicitly representing the complex interconnectivity of the lymphoid tissue network, while treating the dynamics within each lymph node as well–mixed [92]. Cardozo et al. [93, 94] and Jagarapu et al. [95], by contrast, applied a patch model at the within–lymph node scale. Their model considered distinct lymphoid follicles, interconnected via a free blood and lymph compartment, where the spherical lymphoid follicles were in turn subdivided into radial shell patches [95, 93, 94]. Recent experimental and modelling results have lent credibility to this “patch” representation of secondary lymphoid tissues by identifying clear distinctions in HIV viral load and turnover in follicular and extrafollicular compartments, which may also influence viral evolutionary dynamics [96, 97].